OpenStudy (anonymous):

find the limit: lim of x is 9, radical x minus 3/x-9, i have the answer but i need someone to work it out please

OpenStudy (anonymous):

\[\lim_{x\rightarrow 9}\frac{\sqrt{x}-3}{x-9}\]??

OpenStudy (anonymous):

multiply by \[\frac{\sqrt{x}+3}{\sqrt{x}+3}\] to get \[\frac{x-9}{(x-9)(\sqrt{x}-3)}\] then cancel and replace x by 9